# EEG Foundation Challenge: From Cross-Task to Cross-Subject EEG Decoding

### v1

@@ 1.5 Baselines, code, and material provided
The starter kit can be found at: https://eeg2025.github.io/baseline/. This code relies on the Braindecode and EEGDash libraries. These libraries allow data search, loading, fetching, and readily applying deep learning methods on EEG data. In particular, Braindecode bridges the MNE-Python [25] and PyTorch [39] libraries (respectively made for brain signal processing and deep learning). EEGDash allows to train Braindecode models or any PyTorch model from data retrieved from OpenNeuro and NEMAR. In the Braindecode library, we provide a PyTorch model zoo for EEG decoding featuring over 30 deep learning models—including ShallowConvNet [47], EEGNet [33], EEG-Inception [45], EEG-Conformer [50], EEGNex [13], ATCNet [2], BIOT [55], Labram [29] and more, all rigorously standardized and tested for correctness. The provided scripts are only for reference, and the contestants can freely use their codebase for development.
@@ 2.4 Competition promotion and incentives
  • The top three teams will also be recognized as consortium authors on the subsequent NeurIPS competition publication, following the format of our previous NeurIPS competition [54].
To actively encourage newcomers and researchers from under-represented backgrounds, we will offer a dedicated prize of USD 1,000 and NeurIPS workshop registration to the best-performing first-time team led and with the majority of women or members of minority groups, adhering to NeurIPS DIA guidelines. We are committed to inclusivity and will partner with advocacy groups such as Women in AI, Black in AI, Queer in AI, and LatinX in AI to broaden our outreach. Finally, hosting a forum will foster continued engagement beyond the competition, promoting code sharing, collaboration and discussions on EEG decoding to foster collaboration and reproducible research [11].

### v2

@@ 1.5 Baselines, code, and material provided
The starter kit can be found at: https://eeg2025.github.io/baseline/. This code relies on the Braindecode and EEGDash libraries. These libraries allow data search, loading, fetching, and readily applying deep learning methods on EEG data. In particular, Braindecode bridges the MNE-Python Gramfort et al. (2014) and PyTorch Paszke et al. (2019) libraries (respectively made for brain signal processing and deep learning). EEGDash allows to train Braindecode models or any PyTorch model from data retrieved from OpenNeuro and NEMAR. In the Braindecode library, we provide a PyTorch model zoo for EEG decoding featuring over 30 deep learning models—including ShallowConvNet Schirrmeister et al. (2017), EEGNet Lawhern et al. (2018), EEG-Inception Santamaria-Vazquez et al. (2020), EEG-Conformer Song et al. (2022), EEGNex Chen et al. (2024), ATCNet Altaheri et al. (2022), BIOT Yang et al. (2023), Labram Jiang et al. (2024) and more, all rigorously standardized and tested for correctness. The provided scripts are only for reference, and the contestants can freely use their codebase for development.
@@ 2.4 Competition promotion and incentives
  • The top three teams will also be recognized as consortium authors on the subsequent NeurIPS competition publication, following the format of our previous NeurIPS competition Wei et al. (2022).
To actively encourage newcomers and researchers from under-represented backgrounds, we will offer a dedicated prize of USD 1,000 and NeurIPS workshop registration to the best-performing first-time team led and with the majority of women or members of minority groups, adhering to NeurIPS DIA guidelines. We are committed to inclusivity and will partner with advocacy groups such as Women in AI, Black in AI, Queer in AI, and LatinX in AI to broaden our outreach. Finally, hosting a forum will foster continued engagement beyond the competition, promoting code sharing, collaboration and discussions on EEG decoding to foster collaboration and reproducible research Carlson et al. (2025).